OpenStudy (anonymous):

write an equation in point- slope form of the line through point j(4,1) with slope -4

OpenStudy (matlee):

y=-4x+17

OpenStudy (anonymous):

y-1=-4(x-4)

OpenStudy (ahsome):

Equation: \[y – y^1 = m(x – x^1)\] Where: \(y^1\) is \(y\) value for one point \(x^1\) is \(x\) value for one point \(m\) is slope Substitute your values: \[y – y^1 = m(x – x^1)\] \[y – 1 = -4(x – 4)\]

OpenStudy (ahsome):

Hope that helps @UnderCover22

OpenStudy (anonymous):

Thank You Guys

OpenStudy (ahsome):

No problem @UnderCover22 :D. Juse make sure to press the Best Response button to who you think helped you the most :)